### 06 Apr 2026 v1

@@ Abstract
Whether polynomial iteration complexity can be established for smoothing Newton methods (SNMs) in symmetric cone programming (SCP) remains a long-standing open problem. A key difficulty lies in the lack of an analogue of the self-concordant convex framework in interior-point methods (IPMs). In this paper, we answer this question affirmatively. We introduce a reduced smoothing barrier augmented Lagrangian (SBAL) function and prove that it is self-concordant convex-concave, which extends the classical self-concordant theory beyond the convex setting. Furthermore, we show that the parameterized smooth equations associated with SNMs are equivalent to the first-order optimality conditions of a minimax problem whose objective is the reduced SBAL function. Motivated by this equivalence, we propose a path-following smoothing Newton method (PFSNM). The reduced SBAL function induces a central path and an associated neighborhood, which provide estimates of the Newton decrement needed for the path-following analysis. As a result, the method is proven to achieve an iteration complexity of $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$, matching the best-known short-step bound for IPMs. Numerical results on standard benchmarks show that PFSNM is competitive with several well-known interior-point solvers, providing computational support for the polynomial iteration complexity.
@@ 1 Introduction
Symmetric cone programming (SCP) is a fundamental class of convex optimization problems, including linear programming (LP), second-order cone programming (SOCP), semidefinite programming (SDP), and their Cartesian products. Let $\mathbb{E}$ be a Euclidean Jordan algebra equipped with a bilinear operation $\circ$” and an identity element $e$, and let $\mathbb{K}\subseteq\mathbb{E}$ be the associated symmetric cone, i.e., a closed convex cone that is both self-dual and homogeneous. We consider the standard primal-dual form of SCP:
@@
where $c,\,x,\,s\in\mathbb{E}$ and $b,\,\lambda\in\mathbb{R}^{m}$. The linear operator $\mathcal{A}:\mathbb{E}\rightarrow\mathbb{R}^{m}$ is assumed to be surjective, and $\mathcal{A}^{*}$ denotes its adjoint.
Interior-point methods (IPMs) are a standard class of algorithms for symmetric cone programming. They replace the complementarity condition in the Karush–Kuhn–Tucker (KKT) system of (1) by a perturbed relation
$\mathcal{A}x=b,\;\mathcal{A}^{*}\lambda+s=c,\;x,s\in\operatorname{int}\,(\mathbb{K}),\;x\circ s=\mu e,$ (2)
and trace the resulting central path as $\mu\downarrow 0$. Their complexity theory is based on a self-concordant convex framework, which induces a local metric, yields estimates for the Newton decrement, and provides a natural way to define neighborhoods of the central path. In SCP, this framework leads to the classical polynomial iteration bounds: the short-step bound is of order $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$ de Klerk (2002); de Klerk and Vallentin (2016); Schmieta and Alizadeh (2003); Vavasis and Ye (1996); Wright (1997) and the long-step bound is of order $\mathcal{O}(\nu\ln(1/\varepsilon))$ de Klerk (2002); Nesterov (1997); Nocedal and Wright (2006); Schmieta and Alizadeh (2003); Wright (1997), where $\nu$ denotes the rank of $\mathbb{K}$ and $\varepsilon$ is the target accuracy. This worst-case complexity explains why IPMs admit a remarkably robust global theory while retaining strong practical performance.
Alongside IPMs, smoothing Newton methods (SNMs) constitute another important class of algorithms for SCP problems. The basic idea is to reformulate the KKT conditions as a nonsmooth system and then replace the nonsmooth complementarity relation by a parameterized smooth equation
@@
Despite these appealing convergence results, SNMs have long lacked the polynomial iteration complexity guarantee, which remains a long-standing open problem Burke and Xu (1998); Kanzow (1996). A key difficulty is that classical SNMs lack an analogue of the self-concordant convex framework that underlies the polynomial complexity theory of IPMs. The framework provides the central path, its neighborhood structure defined by the merit function, and the Newton-decrement estimates required for path-following analysis. These are fundamental, as they give exact estimates for the descent of the merit function when $\mu$ changes, which are typically hard to obtain without the framework.
@@
Several attempts have been made to address this problem. One direction is to integrate parameterized smooth equations into an interior-point framework. A representative example is the interior-point path-following algorithm proposed by Xu and Burke Xu and Burke (1999), which uses the smoothing CHKS function to generate a rescaled Newton direction within the interior-point framework and achieves polynomial bounds. While important, the iterates are still required to stay in the interior of the cone. In contrast, Hotta, Inaba, and Yoshise Hotta et al. (2000) proposed an SNM based on the smoothing CHKS function that eliminates the interior-point requirement, but at the cost of a non-polynomial complexity of $\mathcal{O}\left({\varepsilon^{-6}}\ln{\varepsilon^{-2}}\right)$, which is far from the standard polynomial complexity bounds of IPMs. Hence, the central question is still open:
Can SNMs for SCP attain the polynomial iteration complexity known for IPMs?
@@
This paper answers this question affirmatively. Our starting point is a reduced smoothing barrier augmented Lagrangian (SBAL) function, which reveals the minimax structure hidden in SNMs. We prove that the function is self-concordant convex-concave — a property that extends the classical self-concordant convex framework to saddle-point problems Nemirovski (1999) and admits a global Newton theory for minimax optimization analogous to that for convex minimization. A further key observation is that the parameterized smooth equations associated with SNMs are exactly the first-order optimality conditions of a minimax problem whose objective is the reduced SBAL function. This equivalence induces a local metric for SNMs, enabling the definition of a central path and an associated neighborhood analogous to those in the interior-point framework. More importantly, it provides the Newton decrement estimates required to control the path-following process. Motivated by this equivalence, we propose a path-following smoothing Newton method (PFSNM) for SCP and establish a worst-case iteration complexity of $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$, which matches the best-known short-step bound for IPMs on symmetric cones. To the best of our knowledge, this is the first polynomial iteration complexity result for a smoothing Newton method in the general SCP setting.
@@
Although our main focus is theoretical, the resulting method is also computationally attractive. PFSNM admits a Newton system with an explicit Schur complement structure, leading to a more efficient system-formation procedure than in existing SNMs. Furthermore, numerical experiments on standard benchmark problems show that PFSNM is competitive with several well-known interior-point solvers. These results are consistent with the established polynomial-complexity theory.
@@ 1.1 Organization
The remainder of the paper is organized as follows. Section 2 reviews preliminaries on Jordan algebras and self-concordant convex-concave functions. Section 3 introduces the reduced SBAL function, establishes its self-concordant convex-concave property, and characterizes the parameterized smooth equations via an equivalent minimax formulation. Section 4 presents the proposed path-following smoothing Newton method and the associated merit functions. Section 5 analyzes the effect of updating the smoothing parameter on the merit functions and finally derives the polynomial iteration complexity of the proposed method. Numerical results are reported in Section 6, validating the effectiveness of PFSNM. The paper concludes in Section 7.
@@ 1.2 Notation
Let $\nabla f(x)$ be the gradient of $f$ at $x$. Then,
@@
For $g\in C^{k}(\hat{\mathbb{E}}\times\mathbb{E})$, denote by $\nabla_{\hat{x}}g(\hat{x},s)\in\hat{\mathbb{E}}$ the partial gradient of $g$ with respect to $\hat{x}$ at $(\hat{x},s)$, and by $D_{\hat{x}}g(\hat{x},s)$ the corresponding partial derivative of $g$. Let $\mathcal{W},\mathcal{H}:\mathbb{E}\to\mathbb{E}$ be linear operators. Write $\mathcal{H}\succ\mathcal{W}$ if
@@
In particular, $\mathcal{H}\succ 0$ means $\langle h,\mathcal{H}h\rangle>0$ for all nonzero $h\in\mathbb{E}$. The boundary and interior of a cone $\mathbb{K}$} are denoted by ${\rm bd}\,(\mathbb{K})$ and ${\rm int}\,({\mathbb{K}})$, respectively. Additional notations and symbols will be introduced as needed.
2 Preliminaries
This section reviews the fundamental concepts that form the basis for our subsequent analysis and algorithmic development. We begin by introducing Euclidean Jordan algebras, which provide the algebraic foundation for symmetric cones and thus play a central role in symmetric cone optimization. We then revisit and generalize the theory of $\alpha$-self-concordant convex-concave functions.
@@ 2.1 Euclidean Jordan algebras and symmetric cones
A Euclidean Jordan algebra $(\mathbb{E},\circ,\langle\cdot,\cdot\rangle)$ is a finite-dimensional real inner product space equipped with a bilinear mapping $\circ:\mathbb{E}\times\mathbb{E}\to\mathbb{E}$ such that, for all $x,y\in\mathbb{E}$,
$x\circ y=y\circ x,\ \,x^{2}\circ(x\circ y)=x\circ(x^{2}\circ y),\ \,\langle x\circ y,z\rangle=\langle y,x\circ z\rangle,$ (4)
where $x^{2}:=x\circ x$. The algebra possesses a unique identity element $e$, satisfying $x\circ e=x$ for all $x\in\mathbb{E}$.
A crucial property of Euclidean Jordan algebras is the existence of a spectral decomposition for any element, which generalizes the eigenvalue decomposition of a symmetric matrix.
@@
Let $\mathbb{E}$ be a Euclidean Jordan algebra of rank $\nu$. For any element $z\in\mathbb{E}$, there exist pairwise orthogonal primitive idempotents $\{v_{1},\ldots,v_{\nu}\}$ and unique real eigenvalues $\lambda_{1}(z),\ldots,\lambda_{\nu}(z)$ such that
@@
By the spectral decomposition, the determinant of $z$ is defined analogously to that of a real matrix:
$\det(z):=\prod\limits_{i=1}^{\nu}\lambda_{i}(z).$ (6)
@@
This definition allows for operations such as the square root $z^{1/2}$, the inverse $z^{-1}$, and the logarithm $\ln(z)$, provided that $z\in\operatorname{int}\,(\mathbb{K})$.
@@
Nonnegative orthant $\mathbb{R}_{+}^{n}=\{x\in\mathbb{R}^{n}\mid x_{i}\geq 0,\,\forall i\}$ $x\circ y:=\operatorname{Diag}(x)y$ $x=\sum_{i=1}^{n}x_{i}e_{i}$
@@
[…]
2.2 Self-concordant convex-concave functions
In this subsection, we introduce the concept and basic properties of $\alpha$-self-concordant convex-concave functions. They provide the theoretical tools for analyzing Newton’s method for finding saddle points, and extend the notion of self-concordant convex functions.
@@ 2.2 Self-concordant convex-concave functions
    $f$ is a barrier for $\mathbb{X}$, i.e., $f(x^{(k)})\to\infty$ along every sequence of points $x^{(k)}\in\mathbb{X}$ converging to the boundary of $\mathbb{X}$;
@@
If $\alpha=1$, $f$ is called standard self-concordant. An $\alpha$-self-concordant convex function $f$ is said to be nondegenerate if the quadratic form $D^{2}f(x)\succ 0$ for all $x\in\mathbb{X}$.
It is well known that an $\alpha$-self-concordant convex function admits a Dikin’s ellipsoid bound, which characterizes the local geometry induced by its second-order derivative.
@@
Let $\mathbb{E}$ be a Euclidean Jordan algebra and $\mathbb{X}\subset\mathbb{E}$ be an open convex domain. Let $f$ be an $\alpha$-self-concordant convex function on $\mathbb{X}$ and $x,\,\Delta x\in\mathbb{E}$. If $r:=\sqrt{\frac{1}{\alpha}D^{2}f(x)[\Delta x,\Delta x]}<1$, then $x+\Delta x\in\mathbb{X}$ and for all $h_{x}\in\mathbb{E}$,
@@
$\phi(x):=-\ln(\det(x)).$ (10)
It follows from Hauser and Güler (2002) that $\phi\in C^{\infty}(\operatorname{int}\,(\mathbb{K}))$ is a nondegenerate 11-self-concordant convex function on $\operatorname{int}\,(\mathbb{K})$. For every $x\in\operatorname{int}\,(\mathbb{K})$, the second-order derivative $D^{2}\phi(x)\succ 0$. Consequently, the inverse operator $(D^{2}\phi(x))^{-1}:\mathbb{E}\to\mathbb{E}$ is well defined and $(D^{2}\phi(x))^{-1}\succ 0$. In particular, the gradient and second-order derivative of $\phi$ satisfy the following properties:
@@
The subsequent analysis focuses on $\alpha$-self-concordant convex-concave functions. We consider an unconstrained minimax problem whose objective is convex in the minimization variables and concave in the maximization variables. To extend Newton’s method to this setting, it is essential to identify a class of convex-concave functions that exhibit similarly favorable geometric properties of self-concordant convex functions. This motivates the introduction of $\alpha$-self-concordant convex-concave functions, a generalization of the definition proposed in Nemirovski (1999).
@@
If $\alpha=1$, the function is called standard self-concordant convex-concave. An $\alpha$-self-concordant convex-concave function $f$ is called nondegenerate if the quadratic form $S_{f}(w)$ is positive definite for all $w\in\hat{\mathbb{E}}\times\mathbb{E}$.
@@
The following proposition relates nondegenerate $\alpha$-self-concordant convex-concave functions to $\alpha$-self-concordant convex functions.
@@
    For every $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ and $h_{1},\,h_{2},\,h_{3}\in\hat{\mathbb{E}}\times\mathbb{E}$, it holds that
@@
Let $f$ be a nondegenerate $\alpha$-self-concordant convex-concave function on $\hat{\mathbb{E}}\times\mathbb{E}$. For any vector $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$, we define two local norms of $h$ associated with $f$ at $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ by
$\|h\|_{f,w,\alpha}=\sqrt{\frac{1}{\alpha}S_{f}(w)[h,h]},\,\,\|h\|^{*}_{f,w,\alpha}=\sqrt{\frac{1}{\alpha}(S_{f}(w))^{-1}[h,h]}.$ (13)
Since $\alpha$ is intrinsic to $f$, we omit it for brevity and write $\|h\|_{f,w}$ and $\|h\|_{f,w}^{*}$ as shorthand for $\|h\|_{f,w,\alpha}$ and $\|h\|_{f,w,\alpha}^{*}$, respectively.
@@
$\delta(w):=\|\Delta w\|_{f,w},\,\,\xi(w):=\|\nabla f\|^{*}_{f,w},\,\,\theta(w):=\max\limits_{\tilde{s}}f(\hat{x},\tilde{s})-\min\limits_{\tilde{x}}f(\tilde{x},s),$ (14)
@@
Let $\hat{\mathbb{E}}$ and $\mathbb{E}$ be finite-dimensional Euclidean spaces, and let $f:\hat{\mathbb{E}}\times\mathbb{E}\to\mathbb{R}$ be a nondegenerate $\alpha$-self-concordant convex-concave function. For every $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ and $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$, the following conclusions hold:
@@
    $\xi(w^{+})\leq\left(\dfrac{\delta(w)}{1-\delta(w)}\right)^{2}\leq\left(\frac{\xi(w)}{1-\xi(w)}\right)^{2}.$ (17)
    Further, if $\delta(w)\leq\xi(w)\leq 2-\sqrt{3}$, then $\xi(w^{+})\leq\dfrac{\delta(w)}{2}\leq\dfrac{\xi(w)}{2}$.

### 21 May 2026 v2

@@ Abstract
It has long remained open whether smoothing Newton methods (SNMs) for symmetric cone programming (SCP) admit polynomial iteration complexity. A key difficulty lies in the lack of an analogue of the self-concordant convex framework underlying interior-point methods (IPMs). In this paper, inspired by Nemirovski’s self-concordant convex-concave theory, we address this open problem by introducing a reduced barrier augmented Lagrangian (BAL) function. We prove that the reduced BAL function is self-concordant convex-concave and establish that the parameterized smooth system arising in SNMs coincides with the first-order optimality conditions of an associated minimax problem. Motivated by this equivalence, we propose a path-following smoothing Newton method (PFSNM). The reduced BAL function induces a central path and an associated neighborhood, which provide estimates for the Newton decrement needed for the path-following analysis. As a result, the method achieves an iteration complexity of $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$, matching the best-known short-step complexity for IPMs. Numerical results on standard benchmarks show that PFSNM is competitive with several well-known interior-point solvers, and the observed performance is consistent with the theoretical development.
@@ 1 Introduction
Symmetric cone programming (SCP) is a fundamental class of convex optimization problems that includes linear programming (LP), second-order cone programming (SOCP), semidefinite programming (SDP), and their Cartesian products. Let $\mathbb{E}$ be a Euclidean Jordan algebra equipped with a bilinear operation $\circ$” and an identity element $e$, and let $\mathbb{K}\subseteq\mathbb{E}$ be the associated symmetric cone, i.e., a closed convex cone that is both self-dual and homogeneous. We consider the standard primal-dual form of SCP:
@@
where $c,\,x,\,s\in\mathbb{E}$ and $b,\,\lambda\in\mathbb{R}^{m}$. The linear operator $\mathcal{A}:\mathbb{E}\rightarrow\mathbb{R}^{m}$ is assumed to be surjective, and $\mathcal{A}^{*}$ denotes its adjoint. We assume that the Slater condition holds for both (P) and (D), which is a standard assumption in SCP.
Interior-point methods (IPMs) are among the most important algorithms for symmetric cone programming. They replace the complementarity condition $x\circ s=0$ in the Karush–Kuhn–Tucker (KKT) system of (1) by $x\circ s=\mu e$ with $\mu>0$, which yields the perturbed KKT system
$\mathcal{A}x=b,\;\mathcal{A}^{*}\lambda+s=c,\;x,s\in\operatorname{int}(\mathbb{K}),\;x\circ s=\mu e.$ (2)
The solutions of this system form the central path. IPMs trace this path as $\mu\downarrow 0$. Their complexity theory is based on a self-concordant convex framework Nesterov and Nemirovskii (1994), which induces a local metric, yields estimates for the Newton decrement, and provides a natural way to define neighborhoods of the central path. In SCP, this framework leads to the classical polynomial iteration complexities: an $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$ iteration complexity for short-step methods De Klerk (2002); De Klerk and Vallentin (2016); Schmieta and Alizadeh (2003); Vavasis and Ye (1996); Wright (1997) and an $\mathcal{O}(\nu\ln(1/\varepsilon))$ iteration complexity for long-step methods De Klerk (2002); Nesterov (1997); Nocedal and Wright (2006); Schmieta and Alizadeh (2003); Wright (1997), where $\nu$ denotes the rank of $\mathbb{K}$ and $\varepsilon$ is the target accuracy. These worst-case complexities explain why IPMs admit a remarkably robust global theory while retaining strong practical performance.
Alongside IPMs, smoothing Newton methods (SNMs) constitute another important class of algorithms for SCP. The basic idea is to reformulate the KKT system as a system of nonsmooth equations and then smooth the complementarity condition. This yields the parameterized smooth system
@@
Despite these appealing convergence results, SNMs have lacked a polynomial iteration complexity guarantee, which has remained a long-standing open problem (see Burke and Xu Burke and Xu (1998) and Kanzow Kanzow (1996)). A key difficulty is that classical SNMs do not admit an analogue of the self-concordant convex framework that underlies the polynomial complexity theory of IPMs. Such a framework provides a central path, a neighborhood structure defined by an appropriate merit function, and the Newton-decrement estimates required for path-following analysis. These are fundamental because they give estimates for the descent of the merit function when $\mu$ is updated, which are typically hard to obtain without the framework.
@@
Several attempts have been made to address this problem. One line of research incorporates the parameterized smooth system into an interior-point framework. A representative example is the interior-point path-following algorithm proposed by Xu and Burke Xu and Burke (1999), which uses the smoothing CHKS function to generate a rescaled Newton direction within the interior-point framework and achieves a polynomial complexity. However, its iterates are still required to stay in the interior of the cone. In contrast, Hotta, Inaba, and Yoshise Hotta et al. (2000) proposed an SNM based on the smoothing CHKS function that eliminates the interior-point requirement, but with a non-polynomial complexity of $\mathcal{O}\left({\varepsilon^{-6}}\ln{\varepsilon^{-2}}\right)$, which is far from the standard polynomial iteration complexities of IPMs. Hence, the following central problem is still open:
Can smoothing Newton methods for symmetric cone programming attain polynomial iteration complexity?
@@
In this paper, we answer this question affirmatively. Our approach is inspired by the self-concordant convex-concave framework introduced by Nemirovski Nemirovski (1999), which provides a natural viewpoint for extending self-concordant ideas from convex minimization to saddle-point problems. The key step is to introduce a reduced barrier augmented Lagrangian (BAL) function that reveals the minimax structure underlying SNMs. We show that this reduced BAL function is self-concordant convex-concave, and that the parameterized smooth system associated with SNMs coincides with the first-order optimality conditions of a minimax problem with the reduced BAL function as the objective. This equivalence induces a local metric for SNMs, enabling the definition of a central path and an associated neighborhood analogous to those in the interior-point framework. More importantly, it provides the Newton decrement estimates required to control the path-following process. On this basis, we propose a path-following smoothing Newton method (PFSNM) for SCP and establish a worst-case iteration complexity of $\mathcal{O}(\sqrt{\nu}\ln(1/\varepsilon))$, which matches the best-known short-step complexity for IPMs on symmetric cones. To the best of our knowledge, this is the first polynomial iteration complexity result for a smoothing Newton method in the general SCP setting.
@@
Although our main focus is theoretical, the resulting method is also computationally attractive. PFSNM admits a Newton system with an explicit Schur-complement structure, leading to an efficient system-formation procedure. Furthermore, numerical experiments on standard benchmark problems show that PFSNM is competitive with several well-known interior-point solvers. These results are consistent with the established polynomial-complexity theory.
@@ 1.1 Organization
The remainder of the paper is organized as follows. Section 2 reviews preliminaries on Euclidean Jordan algebras and self-concordant convex-concave functions. Section 3 introduces the reduced BAL function, establishes its self-concordant convex-concave property, and characterizes the parameterized smooth system via an equivalent minimax formulation. Section 4 presents the proposed path-following smoothing Newton method and the associated merit functions. Section 5 analyzes the effect of updating the smoothing parameter on the merit functions and derives the polynomial iteration complexity of the proposed method. Numerical results are reported in Section 6 and illustrate the practical effectiveness of PFSNM. The paper concludes in Section 7. Auxiliary technical proofs are deferred to Appendix A.
@@ 1.2 Notation
The gradient of $f$ at $x$ is denoted by $\nabla f(x)$ and is defined by
@@
For $g\in C^{k}(\hat{\mathbb{E}}\times\mathbb{E})$ and $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$, we write $\nabla_{\hat{x}}g(w)$ and $\nabla_{s}g(w)$ for the partial gradients, $D_{\hat{x}}g(w)$ and $D_{s}g(w)$ for the corresponding partial derivatives. Let $\mathcal{W},\,\mathcal{H}:\mathbb{E}\to\mathbb{E}$ be linear operators. We write $\mathcal{H}\succ\mathcal{W}$ if
@@
In particular, $\mathcal{H}\succ 0$ means $\langle h,\mathcal{H}h\rangle>0$ for all nonzero $h\in\mathbb{E}$. The interior of a cone $\mathbb{K}$} is denoted by ${\rm int}({\mathbb{K}})$. Additional notation will be introduced as needed.
2 $\alpha$-self-concordant convex-concave functions
This section reviews the core concepts needed in the subsequent analysis. We begin by introducing Euclidean Jordan algebras, which provide the algebraic foundation for symmetric cones and thus play a central role in symmetric cone optimization. We then revisit and generalize the theory of $\alpha$-self-concordant convex-concave functions.
@@ 2.1 Euclidean Jordan algebras and symmetric cones
A Jordan algebra $(\mathbb{E},\circ)$ is a finite-dimensional vector space over the real field $\mathbb{R}$ with a bilinear mapping $\circ:\mathbb{E}\times\mathbb{E}\to\mathbb{E}$ such that, for all $x,y,z\in\mathbb{E}$,
$x\circ y=y\circ x,\ \,x^{2}\circ(x\circ y)=x\circ(x^{2}\circ y),\ \text{where }x^{2}:=x\circ x.$
The algebra is called Euclidean if there exists an inner product, denoted by $\langle\cdot,\cdot\rangle$, which is associative, i.e., for all $x,y,z\in\mathbb{E}$,
$\langle x\circ y,z\rangle=\langle x,y\circ z\rangle.$
A crucial property of Euclidean Jordan algebras is that every element admits a spectral decomposition, which generalizes the eigenvalue decomposition of a symmetric matrix.
@@
Let $\mathbb{E}$ be a Euclidean Jordan algebra of rank $\nu$. For any $z\in\mathbb{E}$, there exist pairwise orthogonal primitive idempotents $\{v_{1},\ldots,v_{\nu}\}$ and unique real eigenvalues $\lambda_{1}(z),\ldots,\lambda_{\nu}(z)$ such that
@@
By the spectral decomposition, the trace and determinant of $z$ are defined analogously to those of a real matrix:
$\operatorname*{tr}(z):=\sum\limits_{i=1}^{\nu}\lambda_{i}(z),\quad\det(z):=\prod\limits_{i=1}^{\nu}\lambda_{i}(z).$ (5)
@@
This definition allows for operations such as the square root $z^{1/2}$, the inverse $z^{-1}$, and the logarithm $\ln(z)$.
By (Faraut and Korányi, 1994, Proposition III.1.5), the bilinear form $\operatorname*{tr}(x\circ y)$ is symmetric, positive definite, and associative. Thus, $\operatorname*{tr}(x\circ y)$ defines an inner product. In the sequel, we denote this inner product by $\langle x,y\rangle$, i.e.,
$\langle x,y\rangle:=\operatorname*{tr}(x\circ y).$ (7)
@@
Nonnegative orthant $\mathbb{R}_{+}^{n}=\{x\in\mathbb{R}^{n}\mid x_{i}\geq 0,\ \forall\,i=1,\dots,n\}$ $x\circ y:=\operatorname{Diag}(x)y$ $x=\sum_{i=1}^{n}x_{i}e_{i}$
Second-order cone $\mathbb{Q}^{n+1}=\left\{(x_{0};\bar{x})\in\mathbb{R}\times\mathbb{R}^{n}\mid x_{0}\geq\|\bar{x}\|_{2}\right\}$ $x\circ y:=\operatorname{Arw}(x)y$a $x=\lambda_{1}v_{1}+\lambda_{2}v_{2}$b
Positive semidefinite cone $\mathbb{S}_{+}^{n}=\{X\in\mathbb{R}^{n\times n}\mid X\succeq 0\}$ $X\circ Y:=\frac{1}{2}(XY+YX)$ $X=\sum_{i=1}^{n}\lambda_{i}v_{i}v_{i}^{\top}$c
a $\operatorname{Arw}(x):=\begin{pmatrix}x_{0}&\bar{x}^{\top}\\ \bar{x}&x_{0}I_{n\times n}\end{pmatrix}$.
b Let $\tilde{v}\in\mathbb{R}^{n}$ satisfy $\|\tilde{v}\|=1$. Then the eigenvalues of $x$ are $\lambda_{1}=x_{0}+\|\bar{x}\|$ and $\lambda_{2}=x_{0}-\|\bar{x}\|$, and the corresponding primitive idempotents are
@@
c $\{\lambda_{i}\}_{i=1}^{n}$ are the eigenvalues of $X$, and $\{v_{i}v_{i}^{\top}\}_{i=1}^{n}$ are the corresponding primitive idempotents.
2.2 $\alpha$-self-concordant convex-concave functions
In this subsection, we review the definition and basic properties of $\alpha$-self-concordant convex-concave functions. They provide the theoretical tools for analyzing Newton’s method for finding saddle points. We begin with the classical notion of self-concordant convex functions.
@@ 2.2 $\alpha$-self-concordant convex-concave functions
    $f$ is a barrier for $\mathbb{X}$, i.e., $f(x^{(k)})\to\infty$ along every sequence of points $x^{(k)}\in\mathbb{X}$ converging to the boundary of $\mathbb{X}$.
@@
If $\alpha=1$, $f$ is called standard self-concordant. An $\alpha$-self-concordant convex function $f$ is said to be nondegenerate if $D^{2}f(x)\succ 0$ for all $x\in\mathbb{X}$.
It is well known that an $\alpha$-self-concordant convex function satisfies a Dikin ellipsoid bound, which characterizes the local geometry induced by its second-order derivative.
@@
Let $\mathbb{E}$ be a Euclidean Jordan algebra and $\mathbb{X}\subset\mathbb{E}$ be an open convex domain. Let $f$ be an $\alpha$-self-concordant convex function on $\mathbb{X}$, $x\in\mathbb{X}$, and $\Delta x\in\mathbb{E}$. If $r:=\sqrt{\frac{1}{\alpha}D^{2}f(x)[\Delta x,\Delta x]}<1$, then $x+\Delta x\in\mathbb{X}$ and for all $h_{x}\in\mathbb{E}$,
@@
$\phi(x):=-\ln(\det(x)).$
By Hauser and Güler (2002), $\phi\in C^{\infty}(\operatorname{int}(\mathbb{K}))$ is a nondegenerate standard self-concordant convex function on $\operatorname{int}(\mathbb{K})$. For every $x\in\operatorname{int}(\mathbb{K})$, the second-order derivative $D^{2}\phi(x)\succ 0$. Consequently, the inverse operator $(D^{2}\phi(x))^{-1}:\mathbb{E}\to\mathbb{E}$ is well defined and satisfies $(D^{2}\phi(x))^{-1}\succ 0$. In particular, the gradient and second-order derivative of $\phi$ satisfy
@@
The subsequent analysis focuses on $\alpha$-self-concordant convex-concave functions. We consider unconstrained minimax problems whose objective is convex in the minimization variables and concave in the maximization variables. To extend Newton’s method to this setting, it is essential to identify a class of convex-concave functions that exhibits geometric properties similar to those of self-concordant convex functions. This motivates the concept of $\alpha$-self-concordant convex-concave functions, which generalizes the definition in Nemirovski (1999).
@@
If $\alpha=1$, $f$ is called standard self-concordant convex-concave. An $\alpha$-self-concordant convex-concave function $f$ is called nondegenerate if the quadratic form $S_{f}(w)$ is positive definite for all $w\in\hat{\mathbb{E}}\times\mathbb{E}$.
@@
The following proposition relates nondegenerate $\alpha$-self-concordant convex-concave functions to nondegenerate $\alpha$-self-concordant convex functions.
@@
    For every $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ and $h_{1},\,h_{2},\,h_{3}\in\hat{\mathbb{E}}\times\mathbb{E}$,
@@
Let $f$ be a nondegenerate $\alpha$-self-concordant convex-concave function on $\hat{\mathbb{E}}\times\mathbb{E}$. For any $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ and any $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$, we define two local norms of $h$ associated with $f$ at $w$ by
$\|h\|_{f,w,\alpha}=\sqrt{\frac{1}{\alpha}S_{f}(w)[h,h]},\,\,\|h\|^{*}_{f,w,\alpha}=\sqrt{\frac{1}{\alpha}(S_{f}(w))^{-1}[h,h]},$ (12)
where $(S_{f}(w))^{-1}[h,h]:=\langle h_{\hat{x}},(D^{2}_{\hat{x}\hat{x}}f(w))^{-1}h_{\hat{x}}\rangle-\langle h_{s},(D^{2}_{ss}f(w))^{-1}h_{s}\rangle$. Since $\alpha$ is fixed once $f$ is specified, we omit it for brevity and write $\|h\|_{f,w}$ and $\|h\|_{f,w}^{*}$ as shorthand for $\|h\|_{f,w,\alpha}$ and $\|h\|_{f,w,\alpha}^{*}$, respectively.
@@
$\delta(w):=\|\Delta w\|_{f,w},\,\,\xi(w):=\|\nabla f(w)\|^{*}_{f,w},\,\,\theta(w):=\max\limits_{\tilde{s}}f(\hat{x},\tilde{s})-\min\limits_{\tilde{x}}f(\tilde{x},s),$ (13)
@@
Let $\hat{\mathbb{E}}$ and $\mathbb{E}$ be finite-dimensional Euclidean spaces, and let $f:\hat{\mathbb{E}}\times\mathbb{E}\to\mathbb{R}$ be a nondegenerate $\alpha$-self-concordant convex-concave function. For every $w=(\hat{x},s)\in\hat{\mathbb{E}}\times\mathbb{E}$ and every $h=(h_{\hat{x}},h_{s})\in\hat{\mathbb{E}}\times\mathbb{E}$, the following statements hold:
@@
    $\xi(w^{+})\leq\left(\dfrac{\delta(w)}{1-\delta(w)}\right)^{2}.$ (16)
    Furthermore, if $\delta(w)\leq\xi(w)\leq 2-\sqrt{3}$, then $\xi(w^{+})\leq\dfrac{\delta(w)}{2}\leq\dfrac{\xi(w)}{2}$.